OpenStudy (anonymous):

what is the perimeter of triangle ABC with vertices A(-2,9), B(7,-3), and C(-2,-3) in the coordinate plane

OpenStudy (anonymous):

a) 21 units b) 15 units c) 34 units d) 36 units

OpenStudy (anonymous):

Use the distance formula to find the length of each vertex

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

distance between (a,b), (c,d): sqrt( (a-c)^2 + (b-d)^2 )

OpenStudy (anonymous):

for example, distance between A(-2, 9) and B(7, -3): sqrt( (-2-7)^2 + (9-(-3))^2 ) = sqrt( 81 + 144) =sqrt( 225) = 15

OpenStudy (anonymous):

do that for vertices BC and AC, then add all of them together